# Sequential Change Detection with Differential Privacy

### v1

@@ Appendix B Proofs of main theorems
In the following we fix the values of Laplace noise $Z_{1},\dots,Z_{t-1}$ and only take probabilities over the randomness of the newly added noise $Z_{t}$ and threshold noise $W$. For any sequence $X_{(1:t)}$, we have:
$\mathbf{P}_{Z_{t},W}(\tilde{T}=t\mid X_{(1:t)})$ $\overset{(i)}{=}\mathbf{P}_{Z_{t},W}(\max_{1\leq j\leq t-1}{\tilde{S_{j}}}<b+W,\tilde{S}_{t}\geq b+W\mid X_{(1:t)})$
@@
$\overset{(iii)}{=}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}f_{Z_{t}}(z)f_{W}(w)\mathbbm{1}[w\in(\max_{1\leq j\leq t-1}{\tilde{S_{j}}}-b,{S}_{t}+z-b]]\,dwdz,$ (22)
@@
$\hat{z}$ $=z+\max_{1\leq j\leq t-1}\tilde{S^{\prime}_{j}}-\max_{1\leq j\leq n-1}\tilde{S_{j}}+S_{t}-S_{t}^{\prime}:=z+\delta_{0},$
@@
Assume $X_{(1:t)}$ and $X^{\prime}_{(1:t)}$ are different at time $k$ only, i.e., $X_{1}=X^{\prime}_{1},\ldots,X_{k-1}=X^{\prime}_{k-1},X_{k+1}=X^{\prime}_{k+1},\ldots,X_{t}=X^{\prime}_{t}$, but $X_{k}\neq X^{\prime}_{k}$. Then we have $S_{1}=S^{\prime}_{1},\ldots,S_{k-1}=S^{\prime}_{k-1},$ but $S_{k}\neq S^{\prime}_{k}$. Without loss of generality, we assume $S_{k}>S^{\prime}_{k}$. By the definition of CUSUM statistics (3), we then have $S_{k+1}\geq S^{\prime}_{k+1}$, $\ldots$, $S_{t}\geq S^{\prime}_{t}$. Since we fixed $Z_{1},\dots,Z_{t-1}$, we have $\tilde{S}_{1}=\tilde{S}^{\prime}_{1},\ldots,\tilde{S}_{k-1}=\tilde{S}^{\prime}_{k-1};\tilde{S}_{k}>\tilde{S}^{\prime}_{k},\tilde{S}_{k+1}\geq\tilde{S}^{\prime}_{k+1},\ldots,\tilde{S}_{t-1}\geq\tilde{S}^{\prime}_{t-1}$. Therefore, we have $\max_{1\leq j\leq t-1}\tilde{S^{\prime}_{j}}-\max_{1\leq j\leq t-1}\tilde{S_{j}}\leq 0$. Moreover, we have
$|\max_{1\leq j\leq t-1}\tilde{S^{\prime}_{j}}-\max_{1\leq j\leq t-1}\tilde{S_{j}}|\leq\max_{1\leq j\leq t-1}|\tilde{S^{\prime}_{j}}-\tilde{S_{j}}|=\max_{1\leq j\leq t-1}|{S^{\prime}_{j}}-{S_{j}}|\leq\Delta,$
@@
$\overset{(iv)}{=}{\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}f_{Z_{t}}(\hat{z}-\delta_{0})f_{W}(\hat{w}-\delta_{1})\mathbbm{1}[\hat{w}-\delta_{1}\in(\max_{1\leq j\leq t-1}{\tilde{S_{j}}}-b,{S}_{t}+\hat{z}-\delta_{0}-b]]\,d\hat{w}d\hat{z}}$
$\overset{(v)}{\leq}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}e^{\epsilon/2}f_{Z_{t}}(\hat{z})e^{\epsilon/2}f_{W}(\hat{w})\mathbbm{1}[\hat{w}\in(\max_{1\leq j\leq t-1}\tilde{S^{\prime}_{j}}-b,S_{t}^{\prime}+\hat{z}-b]]\,d\hat{w}d\hat{z}$
@@
Here $(iv)$ is by change of variables $\hat{z}=z+\delta_{0}$ and $\hat{w}=w+\delta_{1}$, $(v)$ is due to $|\delta_{0}|\leq\Delta$, $|\delta_{1}|\leq\Delta$, and if $Y\sim\text{Lap}(\frac{2\Delta}{\epsilon})$, then $\frac{f_{Y}(y)}{f_{Y}(y^{\prime})}=e^{-\frac{\epsilon}{2\Delta}(|y|-|y^{\prime}|)}\leq e^{\frac{\epsilon}{2\Delta}|y-y^{\prime}|}$, and $(vi)$ is due to the same derivation as in the derivation of Eq. (B). This concludes the proof that $\tilde{T}$ is $\epsilon$-DP. ∎
Proof of Theorm 2.
@@
$\mathbf{E}_{\infty}[\tilde{T}|W=w]$ $\geq x\mathbf{P}_{\infty}(\tilde{T}\geq x|W=w)$
@@
$\mathbf{E}_{\infty}[\tilde{T}|W=w]$ $\geq\frac{e^{\lambda(b+w)}}{4}(1-\lambda)(1-(\frac{2\Delta}{\epsilon}\lambda)^{2})\geq\frac{e^{\lambda(b+w)}}{4}(1-\lambda)(1-\frac{2\Delta}{\epsilon}\lambda),$
@@
    Taking the logarithm and differentiating with respect to $\lambda$, we find the maximizer of the right-hand side is $\lambda^{*}=\frac{\epsilon}{2\Delta}-\frac{2}{b+w}$ and thus
@@
    $\geq\int_{0}^{\infty}e^{\frac{\epsilon}{2\Delta}(b+w)-2}(\frac{2\Delta}{(b+w)\epsilon})^{2}f_{W}(w)dw$
    $=e^{\frac{\epsilon}{2\Delta}b-2}(\frac{2\Delta}{\epsilon})^{2}\int_{0}^{\infty}e^{\frac{\epsilon}{2\Delta}w}\frac{1}{(b+w)^{2}}\frac{\epsilon}{4\Delta}e^{-\frac{\epsilon}{2\Delta}w}dw$
    $=\frac{\Delta}{\epsilon b}\cdot e^{\frac{\epsilon}{2\Delta}b-2}\geq\frac{e^{\frac{\epsilon}{2\Delta}b-2}}{2b}.$
@@
    Similarly, the maximizer of the right-hand side is $\lambda^{*}=1-2/(b+w)$ and thus
@@
    $\mathbf{E}_{\infty}[\tilde{T}$] $=\int_{-\infty}^{\infty}\mathbf{E}_{\infty}[\tilde{T}|W=w]f_{W}(w)dw\geq\int_{0}^{\infty}e^{b+w-2}\frac{1}{(b+w)^{2}}f_{W}(w)dw$
    $=e^{b-2}\int_{0}^{\infty}e^{w}\frac{1}{(b+w)^{2}}\frac{\epsilon}{4\Delta}e^{-\frac{\epsilon}{2\Delta}w}dw=\frac{\epsilon}{4\Delta}e^{b-2}\int_{0}^{\infty}\frac{1}{(b+w)^{2}}e^{-(\frac{\epsilon}{2\Delta}-1)w}dw$
    $\geq\frac{\epsilon}{4\Delta}e^{b-2}\int_{0}^{1}\frac{1}{(b+w)^{2}}e^{-(\frac{\epsilon}{2\Delta}-1)w}dw\geq\frac{\epsilon}{4\Delta}e^{b-2}\frac{1}{(b+1)^{2}}\int_{0}^{1}e^{-\frac{\epsilon}{2\Delta}w}dw$
    $=\frac{\epsilon}{4\Delta}e^{b-2}\frac{1}{(b+1)^{2}}\frac{2\Delta}{\epsilon}(1-e^{-\frac{\epsilon}{2\Delta}})\geq\frac{1}{2}e^{b-2}\frac{1}{(b+1)^{2}}(1-e^{-1})\geq\frac{1}{4}\frac{e^{b-2}}{(b+1)^{2}}.$
Combining these two cases together completes the proof.
@@
$\mathbf{E}_{\tau}[(\tilde{T}(b)-\tau)^{+}|\mathscr{F}_{\tau}]$ $=\mathbf{E}_{\tau}[(\tilde{T}(b)-\tau)\mathbbm{1}(\tilde{T}(b)>\tau)|\mathscr{F}_{\tau}]$
$=\mathbf{E}_{Z_{1},\ldots,Z_{\tau},W}\mathbf{E}_{\tau}[(\tilde{T}(b)-\tau)\mathbbm{1}(\tilde{T}(b)>\tau)|\mathscr{F}_{\tau},Z_{1},\ldots,Z_{\tau},W]$
$\overset{(i)}{\leq}\mathbf{E}_{Z_{1},\ldots,Z_{\tau},W}\mathbf{E}_{\tau}[(\tilde{T}_{0}(b)-\tau)\mathbbm{1}(\tilde{T}_{0}(b)>\tau)|\mathscr{F}_{\tau},Z_{1},\ldots,Z_{\tau},W]$
$\overset{(ii)}{=}\mathbf{E}_{Z_{1},\ldots,Z_{\tau},W}\mathbf{E}_{\tau}[(\tilde{T}_{0}(b)-\tau)\mathbbm{1}(\tilde{T}_{0}(b)>\tau)|Z_{1},\ldots,Z_{\tau},W]$
$=\mathbf{P}_{Z_{1},\ldots,Z_{\tau},W}(\tilde{T}_{0}(b)>\tau)\cdot\mathbf{E}_{\tau}[(\tilde{T}_{0}(b)-\tau)|\tilde{T}_{0}(b)>\tau]$
@@
The inequality $(i)$ is due to $\tilde{T}_{0}(b)\geq\tilde{T}(b)$; the equality $(ii)$ is true because the stopping time $\tilde{T}_{0}$ does not employ any information from $\mathscr{F}_{\tau}$, consequently it is independent from $\mathscr{F}_{\tau}$. The last equality is true because when $\tilde{T}_{0}(b)>\tau$, statistically this is the same as starting at 0 with the change occurring at 0. ∎
@@
$\nu=\inf\{t\geq 1:U_{t}+Z_{t}\geq\tilde{b}\}$. (23)
Here $Z_{t}\sim\text{Lap}(\frac{2\Delta}{\epsilon})$ is the same Laplace noise as in (7). Obviously we have $U_{t}\leq S_{t}$, thus $U_{t}+Z_{t}\leq S_{t}+Z_{t}=\tilde{S}_{t}$ and $\mathbf{E}_{0}[\tilde{T}(b)|W=w]\leq\mathbf{E}_{0}[\nu]$. In the following, we derive an upper bound for $\mathbf{E}_{0}[\nu]$. Note that for notational simplicity, we omit the conditioning on $W=w$ in the following proofs unless explicitly stated otherwise.
Step 1: We first show $\mathbf{E}_{0}[\nu]<\infty$ for any fixed $b>0$. Note that for any integer $t\geq 1$ and $\lambda\in(0,\frac{\epsilon}{2\sqrt{2}\Delta}]$ we have
$\mathbf{P}_{0}(\nu\geq t+1)$ $\leq\mathbf{P}_{0}(U_{t}+Z_{t}<\tilde{b})\leq\frac{\mathbf{E}_{0}[e^{-\lambda(U_{t}+Z_{t})}]}{e^{-\lambda\tilde{b}}}$
@@
$\mathbf{P}_{0}(\nu\geq t+1)$ $\leq 2e^{\lambda\tilde{b}}e^{-t(\lambda I_{0}-\frac{\lambda^{2}\Delta^{2}}{8})}.$
Now take $\lambda=\min\{\frac{4I_{0}}{\Delta^{2}},\frac{\epsilon}{2\sqrt{2}\Delta}\}$, we have $C_{1}:=\lambda I_{0}-\frac{\lambda^{2}\Delta^{2}}{8}>0$, and $\mathbf{P}_{0}(\nu\geq t+1)\leq 2e^{\lambda\tilde{b}}e^{-tC_{1}}$. Thus we have $\mathbf{E}_{0}[\nu]=\sum_{t=1}^{\infty}\mathbf{P}(\nu\geq t)<\infty$.
Step 2: Since $\ell(X_{1}),\ell(X_{2}),\ldots$ are i.i.d. with $\mathbf{E}_{0}[\ell(X_{1})]=I_{0}$, $U_{t}=\sum_{i=1}^{t}\ell(X_{i})$, and $\nu$ is a stopping time with $\mathbf{E}_{0}[\nu]<\infty$, by Wald’s equation [32, Theorem 4.8.6], we have $\mathbf{E}_{0}[U_{\nu}]=\mathbf{E}_{0}[\nu]\mathbf{E}_{0}[\ell(X_{1})]=\mathbf{E}_{0}[\nu]I_{0}$. We rewrite it as follows,
$\mathbf{E}_{0}[\nu]=\frac{\mathbf{E}_{0}[U_{\nu}]}{I_{0}}=\frac{\mathbf{E}_{0}[U_{\nu}+Z_{\nu}-\tilde{b}]+\mathbf{E}_{0}[\tilde{b}-Z_{\nu}]}{I_{0}}=\frac{\tilde{b}+\mathbf{E}_{0}[U_{\nu}+Z_{\nu}-\tilde{b}]+\mathbf{E}_{0}[-Z_{\nu}]}{I_{0}},$ (24)
@@
$\mathbf{E}_{0}[\nu]\leq\frac{\tilde{b}+\Delta+\mathbf{E}_{0}[Z_{\nu}-Z_{\nu-1}]+\mathbf{E}_{0}[-Z_{\nu}]}{I_{0}}=\frac{\tilde{b}+\Delta+\mathbf{E}_{0}[-Z_{\nu-1}]}{I_{0}}\leq\frac{\tilde{b}+\Delta+\mathbf{E}_{0}[|Z_{\nu-1}|]}{I_{0}}.$ (25)
Step 3: We now derive an upper bound for $\mathbf{E}_{0}[|Z_{\nu-1}|]$. Let $m_{0}=\lfloor\frac{\tilde{b}}{2I_{0}}\rfloor-1$ and $m_{1}=\lfloor\frac{2\tilde{b}}{I_{0}}\rfloor+1$. We have
$\mathbf{E}_{0}[|Z_{\nu-1}|]$ $=\sum_{i=1}^{\infty}\mathbf{E}_{0}|Z_{i}\mathbbm{1}(\nu=i+1)|$
[…]
$\hskip 10.0pt+\underbrace{\sum_{i=m_{1}+1}^{\infty}\mathbf{E}_{0}\left(|Z_{i}|\mathbbm{1}(\sum_{j=1}^{i}\ell(X_{j})+Z_{i}<\tilde{b})\right)}_{\text{Part III}}.$
For Part I, by Cauchy-Schwarz inequality, and for any $\lambda\in(0,\frac{\epsilon}{2\sqrt{2}\Delta}]$, we have for any $i\leq m_{0}$,
[…]
$\leq\left(\mathbf{E}_{0}|Z_{i-1}|^{2}\right)^{1/2}\sqrt{\mathbf{P}_{0}(\sum_{j=1}^{i}\ell(X_{j})+Z_{i}\geq\tilde{b})}$
$\leq 2\sqrt{2}(\frac{\Delta}{\epsilon})\sqrt{\frac{\mathbf{E}_{0}e^{\lambda\left(\sum_{j=1}^{i}\ell(X_{j})+Z_{i}\right)}}{e^{\lambda\tilde{b}}}}$
$\leq 2\sqrt{2}(\frac{\Delta}{\epsilon})\sqrt{\frac{e^{(\frac{\lambda^{2}\Delta^{2}}{8}+I_{0}\lambda)i-\lambda\tilde{b}}}{1-4\Delta^{2}\lambda^{2}/\epsilon^{2}}}\leq 4(\frac{\Delta}{\epsilon})e^{-\frac{\lambda}{2}\tilde{b}}e^{\frac{1}{2}(I_{0}\lambda+\frac{\lambda^{2}\Delta^{2}}{8})i},$
where the last inequality is due to $1-4\Delta^{2}\lambda^{2}/\epsilon^{2}\geq\frac{1}{2}$ when $\lambda\leq\frac{\epsilon}{2\sqrt{2}\Delta}$. Thus,
Part I $\leq\frac{4\Delta}{\epsilon}e^{-\frac{\lambda}{2}\tilde{b}}\cdot\sum_{i=1}^{m_{0}}e^{\frac{1}{2}(I_{0}\lambda+\frac{\lambda^{2}\Delta^{2}}{8})i}\overset{(1)}{\leq}\frac{4\Delta}{\epsilon}e^{-\frac{\lambda}{2}\tilde{b}}\cdot\frac{e^{\frac{1}{2}(I_{0}\lambda+\frac{\lambda^{2}\Delta^{2}}{8})(m_{0}+1)}}{\frac{1}{2}(I_{0}\lambda+\frac{\lambda^{2}\Delta^{2}}{8})}\overset{(2)}{\leq}\frac{8\Delta}{\epsilon(I_{0}\lambda+\frac{\lambda^{2}\Delta^{2}}{8})}e^{-\frac{\tilde{b}}{4}\lambda(1-\frac{\lambda\Delta^{2}}{8I_{0}})},$
where $(1)$ is due to $\sum_{i=1}^{m_{0}}e^{Ci}=\frac{e^{C}(e^{Cm_{0}}-1)}{e^{C}-1}\leq\frac{e^{C(m_{0}+1)}}{C}$ for any $C>0$, and $(2)$ is due to $m_{0}+1\leq\tilde{b}/(2I_{0})$. Taking $\lambda=\bar{\lambda}:=\min\{4I_{0}/\Delta^{2},\epsilon/(2\sqrt{2}\Delta)\}$ guarantees $(1-\frac{\bar{\lambda}\Delta^{2}}{8I_{0}})>0$ and thus $\text{Part I}\leq\frac{8\Delta}{\epsilon(I_{0}\bar{\lambda}+\frac{{\bar{\lambda}}^{2}\Delta^{2}}{8})}$.
Similarly for Part III, by Cauchy-Schwarz inequality, and for any $\lambda\in(0,\frac{\epsilon}{2\sqrt{2}\Delta}]$, we have

### v2

@@ Appendix B Proofs of main theorems
In the following, we fix the values of Laplace noise $Z_{1},\dots,Z_{t-1}$ and only take probabilities over the randomness of the newly added noise $Z_{t}$ and threshold noise $W$. For any sequence $X_{(1:t)}$, we have:
$\mathbf{P}_{Z_{t},W}(\tilde{T}=t\mid X_{(1:t)})$ (22)
$\overset{(i)}{=}\mathbf{P}_{Z_{t},W}(\max_{1\leq j\leq t-1}{\tilde{S_{j}}}<b+W,\tilde{S}_{t}\geq b+W\mid X_{(1:t)})$
@@
$\overset{(iii)}{=}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}f_{Z_{t}}(z)f_{W}(w)$
$\hskip 50.0pt\cdot\mathbbm{1}[w\in(\max_{1\leq j\leq t-1}{\tilde{S_{j}}}-b,{S}_{t}+z-b]]\,dwdz,$
@@
$\hat{z}$ $=z+\max_{1\leq j\leq t-1}\tilde{S^{\prime}_{j}}-\max_{1\leq j\leq t-1}\tilde{S_{j}}+S_{t}-S_{t}^{\prime}:=z+\delta_{0},$
@@
Assume $X_{(1:t)}$ and $X^{\prime}_{(1:t)}$ are different at time $k$ only, i.e., $X_{1}=X^{\prime}_{1},\ldots,X_{k-1}=X^{\prime}_{k-1},X_{k+1}=X^{\prime}_{k+1},\ldots,X_{t}=X^{\prime}_{t}$, but $X_{k}\neq X^{\prime}_{k}$. Then we have $S_{1}=S^{\prime}_{1},\ldots,S_{k-1}=S^{\prime}_{k-1},$ but $S_{k}\neq S^{\prime}_{k}$. Without loss of generality, we assume $S_{k}>S^{\prime}_{k}$. By the definition of CUSUM statistics (3), we then have $S_{k+1}\geq S^{\prime}_{k+1}$, $\ldots$, $S_{t}\geq S^{\prime}_{t}$. Since we fixed $Z_{1},\dots,Z_{t-1}$, we have $\tilde{S}_{1}=\tilde{S}^{\prime}_{1},\ldots,\tilde{S}_{k-1}=\tilde{S}^{\prime}_{k-1};\tilde{S}_{k}>\tilde{S}^{\prime}_{k},\tilde{S}_{k+1}\geq\tilde{S}^{\prime}_{k+1},\ldots,\tilde{S}_{t-1}\geq\tilde{S}^{\prime}_{t-1}$. Therefore, we have $\max_{1\leq j\leq t-1}\tilde{S^{\prime}_{j}}-\max_{1\leq j\leq t-1}\tilde{S_{j}}\leq 0$. Moreover, since $\tilde{S}_{k}\geq\tilde{S}^{\prime}_{k}$ for $k=1,\ldots,t-1$, we have
[…]
$=\max_{1\leq j\leq t-1}|{S^{\prime}_{j}}-{S_{j}}|\leq\Delta,$
@@
$\overset{(iv)}{=}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}f_{Z_{t}}(\hat{z}-\delta_{0})f_{W}(\hat{w}-\delta_{1})$
$\hskip 30.0pt\cdot\mathbbm{1}[\hat{w}-\delta_{1}\in(\max_{1\leq j\leq t-1}{\tilde{S_{j}}}-b,{S}_{t}+\hat{z}-\delta_{0}-b]]\,d\hat{w}d\hat{z}$
[…]
$\hskip 70.0pt\cdot\mathbbm{1}[\hat{w}\in(\max_{1\leq j\leq t-1}\tilde{S^{\prime}_{j}}-b,S_{t}^{\prime}+\hat{z}-b]]\,d\hat{w}d\hat{z}$
@@
Here $(iv)$ is by change of variables $\hat{z}=z+\delta_{0}$ and $\hat{w}=w+\delta_{1}$, $(v)$ is due to $|\delta_{0}|\leq\Delta$, $|\delta_{1}|\leq\Delta$, and if $Y\sim\text{Lap}(\frac{2\Delta}{\epsilon})$, then $\frac{f_{Y}(y)}{f_{Y}(y^{\prime})}=e^{-\frac{\epsilon}{2\Delta}(|y|-|y^{\prime}|)}\leq e^{\frac{\epsilon}{2\Delta}|y-y^{\prime}|}$, and $(vi)$ is due to the same derivation as in the derivation of Eq. (22). This concludes the proof that $\tilde{T}$ is $\epsilon$-DP. ∎
Proof of Theorem 2.
@@
$\mathbf{E}_{\infty}[\tilde{T}|W=w]$
$\geq x\mathbf{P}_{\infty}(\tilde{T}\geq x|W=w)$
@@
$\mathbf{E}_{\infty}[\tilde{T}|W=w]$ $\geq\frac{e^{\lambda(b+w)}}{4}(1-\lambda)(1-(\frac{2\Delta}{\epsilon}\lambda)^{2})$
$\geq\frac{e^{\lambda(b+w)}}{4}(1-\lambda)(1-\frac{2\Delta}{\epsilon}\lambda),$
@@
    Taking the logarithm and differentiating with respect to $\lambda$, we find the unique stationary point of the right-hand side is $\lambda^{*}=\frac{\epsilon}{2\Delta}-\frac{2}{b+w}.$ Moreover, if $w>4\Delta/\epsilon,$ we have $\lambda^{*}\in(0,\frac{\epsilon}{2\Delta}),$ which implies $\lambda^{*}$ is the maximizer of the right-hand side. Thus
@@
    $\geq\int_{4\Delta/\epsilon}^{\infty}e^{\frac{\epsilon}{2\Delta}(b+w)-2}(\frac{2\Delta}{(b+w)\epsilon})^{2}f_{W}(w)dw$
    $=e^{\frac{\epsilon}{2\Delta}b-2}(\frac{2\Delta}{\epsilon})^{2}\int_{{4\Delta/\epsilon}}^{\infty}e^{\frac{\epsilon}{2\Delta}w}\frac{1}{(b+w)^{2}}\frac{\epsilon}{4\Delta}e^{-\frac{\epsilon}{2\Delta}w}dw$
    $=\frac{\Delta}{\epsilon(b+4\Delta/\epsilon)}\cdot e^{\frac{\epsilon}{2\Delta}b-2}\geq\frac{e^{\frac{\epsilon}{2\Delta}b-2}}{2b+4}\geq\frac{e^{\frac{\epsilon}{2\Delta}b-2}}{4(b+1)^{2}}.$
@@
    Similarly, the unique stationary point of the right-hand side is $\lambda^{*}=1-2/(b+w).$ Moreover, if $b>2,$ we have $\lambda^{*}\in(0,\frac{\epsilon}{2\Delta}),$ which implies that $\lambda^{*}=1-2/(b+w)$ is also the maximizer of the right-hand side. Thus,
@@
    $\geq\int_{0}^{\infty}e^{b+w-2}\frac{1}{(b+w)^{2}}f_{W}(w)dw$
    $=e^{b-2}\int_{0}^{\infty}e^{w}\frac{1}{(b+w)^{2}}\frac{\epsilon}{4\Delta}e^{-\frac{\epsilon}{2\Delta}w}dw$
    $\geq\frac{\epsilon}{4\Delta}e^{b-2}\int_{0}^{\infty}\frac{1}{(b+w)^{2}}e^{-\frac{\epsilon}{2\Delta}w}dw$
    $\geq\frac{\epsilon}{4\Delta}e^{b-2}\int_{0}^{1}\frac{1}{(b+w)^{2}}e^{-\frac{\epsilon}{2\Delta}w}dw$
    $\geq\frac{\epsilon}{4\Delta}e^{b-2}\frac{1}{(b+1)^{2}}\int_{0}^{1}e^{-\frac{\epsilon}{2\Delta}w}dw$
    $=\frac{\epsilon}{4\Delta}e^{b-2}\frac{1}{(b+1)^{2}}\frac{2\Delta}{\epsilon}(1-e^{-\frac{\epsilon}{2\Delta}})$
    $\geq\frac{1}{2}e^{b-2}\frac{1}{(b+1)^{2}}(1-e^{-1})\geq\frac{1}{4}\frac{e^{b-2}}{(b+1)^{2}}.$
Combining these two cases together completes the proof. ∎
@@
$\mathbf{E}_{\tau}[(\tilde{T}(b)-\tau)^{+}|\mathscr{F}_{\tau}]$
$=\mathbf{E}_{\tau}[(\tilde{T}(b)-\tau)\mathbbm{1}(\tilde{T}(b)>\tau)|\mathscr{F}_{\tau}]$
$=\mathbf{E}_{Z_{1},\ldots,Z_{\tau},W}\mathbf{E}_{\tau}[(\tilde{T}(b)-\tau)\mathbbm{1}(\tilde{T}(b)>\tau)|\mathscr{F}_{\tau},Z_{1},...,Z_{\tau},W]$
$\overset{(i)}{\leq}\!\mathbf{E}_{Z_{1},...,Z_{\tau},W}\mathbf{E}_{\tau}[(\tilde{T}_{0}(b)\!-\!\tau)\mathbbm{1}(\tilde{T}_{0}(b)\!>\!\tau)|\mathscr{F}_{\tau},Z_{1},...,Z_{\tau},W]$
$\overset{(ii)}{=}\!\mathbf{E}_{Z_{1},\ldots,Z_{\tau},W}\mathbf{E}_{\tau}[(\tilde{T}_{0}(b)-\tau)\mathbbm{1}(\tilde{T}_{0}(b)>\tau)|Z_{1},...,Z_{\tau},W]$
$\overset{(iii)}{=}\mathbf{P}_{Z_{1},\ldots,Z_{\tau},W}(\tilde{T}_{0}(b)>\tau)\cdot\mathbf{E}_{\tau}[(\tilde{T}_{0}(b)-\tau)|\tilde{T}_{0}(b)>\tau]$
@@
The inequality $(i)$ is due to $\tilde{T}_{0}(b)\geq\tilde{T}(b)$; the equality $(ii)$ is true because the stopping time $\tilde{T}_{0}$ does not employ any information from $\mathscr{F}_{\tau}$, consequently it is independent from $\mathscr{F}_{\tau}$; the equality $(iii)$ is due to the fact that $\mathbbm{1}(\tilde{T}_{0}(b)>\tau)\in\sigma\{Z_{1},\cdots,Z_{\tau},W\}.$ The last equality is true because when $\tilde{T}_{0}(b)>\tau$, statistically this is the same as starting at 0 with the change occurring at 0. ∎
@@
$\nu(\tilde{b})=\inf\{t\geq 1:U_{t}+Z_{t}\geq\tilde{b}\}$. (23)
[…]
$\mathbf{P}_{0}(\nu(\tilde{b})\geq t+1)$ $\leq\mathbf{P}_{0}(U_{t}+Z_{t}<\tilde{b})\leq\frac{\mathbf{E}_{0}[e^{-\lambda(U_{t}+Z_{t})}]}{e^{-\lambda\tilde{b}}}$
@@
$\mathbf{P}_{0}(\nu(\tilde{b})\geq t+1)$ $\leq 2e^{\lambda\tilde{b}}e^{-t(\lambda I_{0}-\frac{\lambda^{2}\Delta^{2}}{8})}.$
Now take $\lambda=\min\{\frac{4I_{0}}{\Delta^{2}},\frac{\epsilon}{2\sqrt{2}\Delta}\}$, we have $C_{1}:=\lambda I_{0}-\frac{\lambda^{2}\Delta^{2}}{8}>0$, and $\mathbf{P}_{0}(\nu(\tilde{b})\geq t+1)\leq 2e^{\lambda\tilde{b}}e^{-tC_{1}}$. Thus we have $\mathbf{E}_{0}[\nu(\tilde{b})]=\sum_{t=1}^{\infty}\mathbf{P}(\nu(\tilde{b})\geq t)<\infty$.
[…]
$\mathbf{E}_{0}[\nu]$ $=\frac{\mathbf{E}_{0}[U_{\nu}]}{I_{0}}=\frac{\mathbf{E}_{0}[U_{\nu}+Z_{\nu}-\tilde{b}]+\mathbf{E}_{0}[\tilde{b}-Z_{\nu}]}{I_{0}}$ (24)
$=\frac{\tilde{b}+\mathbf{E}_{0}[U_{\nu}+Z_{\nu}-\tilde{b}]+\mathbf{E}_{0}[-Z_{\nu}]}{I_{0}},$
@@
$\mathbf{E}_{0}[\nu]$ $\leq\frac{\tilde{b}+\Delta+\mathbf{E}_{0}[Z_{\nu}-Z_{\nu-1}]+\mathbf{E}_{0}[-Z_{\nu}]}{I_{0}}$ (25)
$=\frac{\tilde{b}+\Delta+\mathbf{E}_{0}[-Z_{\nu-1}]}{I_{0}}\leq\frac{\tilde{b}+\Delta+\mathbf{E}_{0}[|Z_{\nu-1}|]}{I_{0}}.$
Step 3: We now derive an upper bound for $\mathbf{E}_{0}[|Z_{\nu-1}|]$. Let $m_{1}=\lfloor\frac{2\tilde{b}}{I_{0}}\rfloor+1$. We have
$\mathbf{E}_{0}[|Z_{\nu-1}|]=\sum_{i=1}^{\infty}\mathbf{E}_{0}|Z_{i}\mathbbm{1}(\nu=i+1)|$ (26)
$\leq\underbrace{\sum_{i=1}^{m_{1}}\mathbf{E}_{0}|Z_{i}\mathbbm{1}(\nu=i+1)|}_{\text{Part I}}$
$\hskip 20.0pt+\underbrace{\sum_{i=m_{1}+1}^{\infty}\mathbf{E}_{0}\left(|Z_{i}|\mathbbm{1}(\sum_{j=1}^{i}\ell(X_{j})+Z_{i}<\tilde{b})\right)}_{\text{Part II}}.$
For Part I, note
Part I $\leq\sum_{i=1}^{m_{1}}(\mathbf{E}_{0}|Z_{i}|^{2})^{1/2}\sqrt{\mathbf{P}_{0}(\nu=i+1)}$
$\leq 2\sqrt{2}(\frac{\Delta}{\epsilon})\left(\sum_{i=1}^{m_{1}}\sqrt{\mathbf{P}_{0}(\nu=i+1)}\right)$
$\leq 2\sqrt{2}(\frac{\Delta}{\epsilon})\sqrt{(\sum_{i=1}^{m_{1}}1)(\sum_{i=1}^{m_{1}}\mathbf{P}_{0}(\nu=i+1))}$
$\leq 2\sqrt{2}(\frac{\Delta}{\epsilon})\sqrt{m_{1}}$
$\leq 2\sqrt{2}(\frac{\Delta}{\epsilon})\sqrt{\frac{2\tilde{b}}{I_{0}}+1}\leq 2\sqrt{2}(\frac{\Delta}{\epsilon})(\sqrt{\frac{2\tilde{b}}{I_{0}}}+1)$
$=4\frac{\Delta}{\epsilon}\sqrt{\frac{\tilde{b}}{I_{0}}}+2\sqrt{2}\frac{\Delta}{\epsilon}.$ (27)
For Part II, by Cauchy-Schwarz inequality, and for any $\lambda\in(0,\frac{\epsilon}{2\sqrt{2}\Delta}]$, we have